origin point and roots of equation z^2+pz+q =0 makes a equilliteral triangle if
  1. p^2=2q
  2. p^2=3q

Grade:11

1 Answers

Vikas TU
7 years ago
let z1 and z2 be the roots then
z1 + z2 = -p
z1*z2 = q
Now Use rotation theorem,
z1 = Me^ia
and
z2 = Me^i(a + pi/3)
z2/z1 = e^ipi/3 = > ½ + i*root3/2.
now u can do.